OpenStudy (anonymous):

Without graphing, tell how many x-intercepts y= 2x2 + 4x + 2 has.

OpenStudy (abb0t):

one, and they are teh same number.

OpenStudy (abb0t):

well, two, but they are the the same number is what I mean.

Parth (parthkohli):

We can let \(y = 0\) and then solve \(2x^2 + 4x + 2 = 0\).

OpenStudy (anonymous):

if you solve the equation, u end up with (x+1)^2, so, the the function will be satisfied by 1, but twice....so it has one x intercept

OpenStudy (blurbendy):

Two intercepts. 2x^2 + 4 x + 2. The leading exponential power tells you how many intercepts will be present (in this case, it is 2) For example, if it were 3x^3 it would have 3 x intercepts

OpenStudy (anonymous):

sorry, its -1

OpenStudy (abb0t):

@tanvirms how did you get -1?

OpenStudy (anonymous):

2x^2+2x+1+0 or, x^2_2x+1=0 or, (x+1)^2=0 or, X+1=0 or, x= -1

OpenStudy (anonymous):

sorry, 2x^2+4x+2=0 *

OpenStudy (anonymous):

thanks guys !